7 over 9 divided 5 over 12 plus parentheses [1 over 2] power of 2.
Answer by CubeyThePenguin(3113)

(7/9)/(5/12) + (1/2)^2
= 84/45 + 1/4
= (84 * 4 + 45)/180
= 381/180